# Lifelong Learning with Behavior Consolidation for Vehicle Routing

### v2

@@ 4.6 Applicability
We further implement LLR-BC on Omni Zhou et al. (2023) and INViT Fang et al. (2024), and compare it with fine-tuning on task order 1. As Table 4 shows, LLR-BC outperforms fine-tuning generally, with a similar pattern found in on POMO. More details of experiment settings, results and discussions are presented in Appendix D.5. In summary, LLR-BC is effective on the focused lifelong learning scenario, without relying on specific characteristics of the base neural solver.
@@ 5 Conclusions
We propose LLR-BC, a lifelong learning framework for neural VRP solvers, considering tasks with varying problem scales and distributions that arrive sequentially and require to be learned. We introduce two core components: CaEW, which emphasizes low-confidence behaviors, and DsBC, which preserves learned behavior effectively in a decision-seeking way. Experiments across diverse orders of tasks with varying distributions and scales demonstrate that LLR-BC consistently outperforms baselines in terms of performance in solving learned tasks, resistance to catastrophic forgetting, learning efficiency on new tasks, and zero-shot generalization. Ablation and sensitivity studies confirm the effectiveness of CaEW and DsBC and the framework’s robustness to hyperparameter settings. LLR-BC is both model- and RL algorithm-agnostic, and can be integrated with various neural VRP solvers. Experiments on different base neural solvers verify LLR-BC’s applicability.
Although studied on cross-distribution and cross-scale settings, LLR-BC can also be applied to other lifelong learning scenarios of neural VRP solvers, such as tasks that differ in their distance matrices or correspond to different problem variants. LLR-BC can further remove the reliance on task identity and adapt to scenarios with continuously changing tasks by applying reservoir sampling during training on each instance instead of only the final epoch of each task. Despite its strong performance, LLR-BC has limitations in some specific cases. A fixed $|\mathcal{E}|$ could be difficult to tune when task scales differ greatly, and reservoir sampling in the last epoch favors tasks with more experiences generated in one epoch, leading to potential buffer imbalance. Future work on addressing these issues could expand its applicability.
@@ A.1 Overall Process of LLR-BC
For a sequence of tasks, LLR-BC conducts lifelong learning with Algorithm 1. For each new task in the lifelong learning process, LLR-BC trains the solver model $\pi_{\theta}$ and updates the experience buffer $\mathcal{B}$ as Algorithm 2 demonstrates. Only the model parameter $\theta$ and the buffer $\mathcal{B}$ are transferred between tasks. LLB-RC can sequentially train on any number of tasks, as it does not rely on any knowledge about the number or order of tasks.
@@ A.3 Reverse KLD
where $P(i)$ and $Q(i)$ are the probability of candidate $i$ of the target distribution $P$ and of another distribution $Q$. KLD is asymmetric, i.e., $D_{KL}(P||Q)\neq D_{KL}(Q||P)$, although both attain their minimum when $P=Q$. By switch $Q$ and $P$ KLD convert to RKLD, i.e.,
@@
When the teacher’s action probability $P(i)$ is close to zero, the KLD for action $i$ becomes very small. As a result, using KLD as a loss term places little emphasis on the mismatch between $P(i)$ and the learner’s probability $Q(i)$ in such cases. This leads to overly mild penalties when the learner assigns high probability to actions that the teacher has effectively ruled out, i.e., learned incorrect actions. Consequently, the learner tends to spread probability mass across all actions to cover possible modes, i.e., mode-covering.
In contrast, RKLD places greater emphasis on penalizing mismatches when the learner’s probability $Q(i)$ is high. This makes RKLD more effective at discouraging the learner from assigning high probability to those learned incorrect actions, thereby encouraging focus on the most probable teacher actions. As a result, RKLD promotes mode-seeking or decision-seeking behavior.
@@
A.4 Computational Cost
@@
Typically, in a training batch of problems with scale $N$, fine-tuning constraints at least $N$ steps of model forward propagation to generate complete solutions. Compared with fine-tuning the base neural solver, LLR-BC introduces additional time cost from computing the new behaviors of the current model on $|\mathcal{E}|$ sampled old states, storing and sampling experiences from the buffer, and calculating the RKLD between new and old behaviors as the LBC. The first part requires $|\mathcal{E}|$ additional forward propagations, while the latter two involve only lightweight arithmetic operations, whose run cost is negligible. Therefore, the extra time cost introduced by LLR-BC per problem batch is $O(|\mathcal{E}|*t_{m})$, $|\mathcal{E}|*t_{m}$, with $t_{m}$ denotes the complexity of forward propagation. With $|\mathcal{E}|=4$ (already outperforming the compared methods) and $N=20,50,100$, ideally, LLR-BC incurs only an additional computational overhead of less than 20%, 8%, and 2%, respectively, compared with fine-tuning. Under our implementation, for all experiments, LLR-BC introduces no more than 8 additional hours based on fine-tuning of training for the whole lifelong learning process. Given its superior performance, this additional training complexity is a reasonable and acceptable trade-off.
A.5 Novelty
LLR-BC aims to mitigate the catastrophic forgetting issue when learning from multiple tasks sequentially, by consolidating the solver’s good behaviors on previous critical states. LLR-BC is novel and has significant differences from existing methods.
LLR-BC differs from existing methods for improving cross-distribution cross-scale zero-shot generalization ability (i.e., Fang et al. (2024); Jiang et al. (2022); Gao et al. (2024)) in that LLR-BC works at a higher level, involving a lifelong learning process on a sequence of tasks. These existing methods can serve as the base neural solver in LLR-BC.
LLR-BC differs from the method proposed by Li et al. (2024); Feng et al. (2025) in that: their methods conduct experience replay at the problem instance level, retrain solvers on new instances actively generated from previous tasks and equally learns from each step of solving one instance. In contrast, LLR-BC relies on only buffering problem-solving steps (state-bahavior pairs) generated in the past instead of generating new problem instances of previous tasks. LLR-BC consolidates knowledge in units of a single step (one state-behavior pair) instead of a whole trajectory of solving an instance, and emphasizes more important steps, making it lower complexity and potentially more efficient.
In addition, this work also differs from  Li et al. (2024); Feng et al. (2025) in terms of the breadth of experimentation and performance evaluation. These two works consider only a single task order and make strong assumptions about the order. Although their methods have general applicability, their experiments are conducted solely with POMO as the base neural solver, and performance evaluation is limited to testing the final model—after learning all tasks—on a subset of previously seen tasks. In contrast, our work does not rely on assumptions about task order; we validate performance across multiple randomized task orders, test several widely studied solvers, and introduce a comprehensive set of metrics to enable broad and systematic performance evaluation.
@@ C.1 Baselines
To adapt LiBOG in VRP tasks, we directly embed its two key modules, i.e., the intra-task and inter-task consolidation terms, into POMO for lifelong learning. After a brief parameter search, the configuration that achieved the best AP performance was selected: the weight of the consolidation term is $10$ for EWC and LiBOG, and the weight of the intra-task consolidation term is 1e41e-4 for LiBOG. For all runs, the solver learns the first task from an identical randomly generated model. As Restart learns each task independently under any task order without being affected by the order, we run Restart once on each task and calculate metrics under different task orders. Therefore, Restart’s APl owns not std. value. As in most cases, POMO can complete the learning on each single task for 200 epochs (Kwon et al., 2020), we let each method train for 200 epochs on each task. To ensure comparable training time across tasks of different problem scales, we vary the number of training problems sampled per epoch: 10,000 for tasks of scale 20, 4,000 for scale 50, and 2,000 for scale 100, with batch sizes of 64, 32, and 16, respectively. Following  Kwon et al. (2020), during training, we sample a batch of problem instances and run $N$ parallel solving processes per instance, each corresponding to a different starting node, where $N$ is the problem scale. This results in $N^{2}$ parallel trajectories per batch, yielding $N^{2}$ independent states and behaviors per experience batch. The size of the experience batch varies across tasks with different problem scales. We consider this reasonable, as larger-scale tasks are typically more complex and thus require larger experience batches to effectively retain learned knowledge. We experimented with buffering and sampling at the level of individual experiences rather than experience batches. However, when using GPU-based parallelization, operating at the batch level significantly reduces computation time without yielding significant performance differences.
@@ C.2 Evaluation Metrics
C.3 Other Details
@@ D.1 Main Experiments on CVRP
Order Method k=3 k=6 AP AF AMF APl AG AP AF AMF APl AG 1 Restart 72.0 8.2 8.2 11.8 36.3 95.7 27.1 27.7 9.1 40.2 Fine-tuning 32.8 3.2 3.2 7.5 32.8 34.5 19.9 23.1 4.6 31.3 EWC 35.9 3.1 3.1 8.6 32.6 39.6 19.7 21.5 7.5 33.5 LiBOG 38.3 2.7 2.7 8.9 33.4 47.7 18.0 18.6 8.3 34.7 LLR-BC 8.0 0.0 0.0 7.0 27.1 4.7 0.6 0.7 4.1 23.3 2 Restart 44.9 92.5 92.5 8.7 60.7 42.1 37.2 64.4 9.1 52.6 Fine-tuning 21.8 29.9 29.9 3.7 50.0 18.8 14.3 23.3 2.5 38.4 EWC 19.5 33.1 33.1 4.4 48.6 19.7 12.4 23.3 4.9 37.9 LiBOG 22.4 29.1 29.1 4.9 51.1 22.2 12.1 21.2 5.4 39.2 LLR-BC 2.8 0.0 0.0 3.3 37.1 2.2 0.3 0.3 1.8 22.6 3 Restart 106.8 10.7 10.7 7.8 30.9 95.7 22.8 35.7 9.1 51.9 Fine-tuning 35.6 4.9 4.9 6.1 24.1 33.4 21 27.4 3.9 44.4 EWC 40.3 4.1 4.1 6.7 22.3 38.0 18.0 25.4 6.8 40.5 LiBOG 42.6 4.3 4.3 6.8 25.0 43.0 18.4 25.9 6.7 41.7 LLR-BC 6.4 0.0 0.0 5.9 23.0 5.0 0.9 1.0 3.7 23.1 4 Restart 35.7 44.5 44.5 10.3 77.3 42.1 26.5 35.5 9.1 52.1 Fine-tuning 14.6 32.6 32.6 7.0 70.0 20.2 18.9 29.9 4.5 45.8 EWC 19.2 28.4 28.4 8.4 69.6 27.1 18.5 23.6 7.9 46.4 LiBOG 19.4 29.1 29.1 8.7 69.0 26.5 20.0 24.3 7.9 46.7 LLR-BC 7.0 0.3 0.3 6.4 58.2 4.7 0.3 0.6 4.0 32.6 5 Restart 27.2 19.3 19.3 9.5 58.6 26.7 93.1 96.5 9.1 50.5 Fine-tuning 11.8 11.8 11.8 7.0 50.2 10.5 25.5 36.7 3.7 50.6 EWC 17.8 14.5 14.5 8.2 49.1 17.0 29.0 33.7 7.5 40.9 LiBOG 18.3 13.5 13.5 8.6 49.0 17.3 29.9 35.7 7.7 41.7 LLR-BC 6.5 0.4 0.4 6.7 45.3 4.2 1.6 1.6 3.7 31.8

### v3

@@ 4.6 Applicability
We further implement LLR-BC on Omni Zhou et al. (2023) and INViT Fang et al. (2024), and compare it with fine-tuning on task order 1. As Table 4 shows, LLR-BC outperforms fine-tuning generally, with a similar pattern found on POMO. More details of experiment settings, results, and discussions are presented in Appendix D.6. In summary, LLR-BC is effective on the focused lifelong learning scenario, without relying on specific characteristics of the base neural solver.
@@ 5 Conclusions
We propose LLR-BC, a lifelong learning framework for neural VRP solvers, considering tasks with varying problem scales and distributions that arrive sequentially and need to be learned. We introduce two core components: CaEW, which emphasizes low-confidence behaviors, and DsBC, which preserves learned behavior effectively in a decision-seeking way. Experiments across diverse orders of tasks with varying distributions and scales demonstrate that LLR-BC consistently outperforms baselines in terms of performance in solving learned tasks, resistance to catastrophic forgetting, learning efficiency on new tasks, and zero-shot generalization. Ablation and sensitivity studies confirm the effectiveness of the core designs of LLR-BC and the framework’s robustness to hyperparameter settings. LLR-BC is both model- and RL algorithm-agnostic, and can be integrated with various neural VRP solvers. Experiments on different base neural solvers verify LLR-BC’s applicability.
Although studied focus on cross-distribution and cross-scale settings, LLR-BC is in principle applicable to other lifelong learning scenarios for neural VRP solvers, such as tasks that differ in their distance matrices. Lifelong learning across different problem variants (e.g., TSP and CVRP) is also a practically important research direction, which we leave for future work, although it may require task-specific model components (Drakulic et al., 2025). LLR-BC could further reduce its reliance on task identity and adapt to scenarios with continuously evolving tasks by applying reservoir sampling during training on each instance, rather than only at the final epoch of each task. Despite its strong performance, LLR-BC has limitations in certain cases. For example, a fixed $|\mathcal{E}|$ can be difficult to tune when task scales differ substantially, potentially leading to reduced plasticity on small-scale tasks, where old experiences dominate new ones in a batch, and reduced stability on large-scale tasks, where new experiences dominate old ones in a batch. Addressing these issues in future work could further broaden the applicability of LLR-BC.
Reproducibility Statement
All algorithmic details (cf. Section 3 and Appendix A), training protocols (cf. Section 4 and Appendix C), and evaluation metrics (cf. Section 4 and Appendix C.2) are described in the main paper and further elaborated in the Appendix. For empirical studies, we provide a detailed description of the datasets and instance generation pipeline (cf. Appendix B). Hyperparameters and implementation details for all baselines are also reported in Section 4 and Appendix C. Upon acceptance, we will release our code and scripts for reproducing our experiments, including instructions for running and data preparation. Together, these resources enable independent researchers to replicate our results and build upon our contributions.
@@ A.1 Overall Process of LLR-BC
For a sequence of tasks, LLR-BC conducts lifelong learning with Algorithm 1. For each new task in the lifelong learning process, LLR-BC trains the solver model $\pi_{\theta}$ and updates the experience buffer $\mathcal{B}$ as Algorithm 2 demonstrates. Only the model parameter $\theta$ and the buffer $\mathcal{B}$ are transferred between tasks. LLR-BC can sequentially train on any number of tasks, as it does not rely on any knowledge about the number or order of tasks.
@@ A.3 Reverse KLD
where $P(i)$ and $Q(i)$ are the probability of candidate $i$ of the target/teacher distribution $P$ and of another distribution $Q$. KLD is asymmetric, i.e., $D_{KL}(P||Q)\neq D_{KL}(Q||P)$, although both attain their minimum when $P=Q$. By switching $Q$ and $P$ KLD convert to RKLD, i.e.,
@@
When the teacher’s action probability $P(i)$ is close to zero, the KLD for action $i$ becomes very small. As a result, using KLD as a loss term places little emphasis on the mismatch between $P(i)$ and the learner’s probability $Q(i)$ in such cases. This leads to overly mild penalties when the learner assigns high probability to actions that the teacher has effectively ruled out, i.e., learned incorrect actions. Consequently, the learner tends to spread probability mass across all actions to cover possible modes (Chan et al., 2022; Wu et al., 2025).
In contrast, RKLD places greater emphasis on penalizing mismatches when the learner’s probability $Q(i)$ is high. This makes RKLD more effective at discouraging the learner from assigning high probability to those learned incorrect actions, thereby encouraging focus on the most probable teacher actions. As a result, RKLD promotes mode-seeking behavior (Chan et al., 2022; Wu et al., 2025).
@@
A.4 Confidence-aware Weighting
A low-confidence decision of the solver can arise in two situations: (i) there is a single best action, but the model has not fully learned the current state and therefore exhibits high uncertainty; or (ii) there are multiple equally good actions. Since we only buffer experiences at the final epoch of each task, when the model is already well trained on that task, both types of low-confidence decisions deserve higher emphasis (larger weights): (i) high uncertainty indicates that the state is difficult to learn and thus crucial for learning the task [3,4]; and (ii) multiple equally good actions, which mostly lead to similarly good solutions, create plateaus in the solution space and make the optimization problem harder to solve.
CaEW is designed to assign smaller weights to steps where the model selects the estimated best action with higher probability (i.e., higher decision confidence). Besides variance, other probability-based confidence measures with similar properties could also be used. CaEW is not specialized for any particular confidence measure. We choose variance because it is simple, straightforward, and effective.
A.5 Computational Cost
@@
Typically, in a training batch of problems with scale $N$, fine-tuning constraints at least $N$ steps of model forward propagation to generate complete solutions. Compared with fine-tuning the base neural solver, LLR-BC introduces additional time cost from computing the new behaviors of the current model on $|\mathcal{E}|$ sampled old states, storing and sampling experiences from the buffer, and calculating the RKLD between new and old behaviors as the LBC. The first part requires $|\mathcal{E}|$ additional forward propagations, while the latter two involve only lightweight arithmetic operations, whose run cost is negligible. Therefore, the extra time cost introduced by LLR-BC per problem batch is $O(|\mathcal{E}|*t_{m})$, $|\mathcal{E}|*t_{m}$, with $t_{m}$ denotes the complexity of forward propagation. With $|\mathcal{E}|=4$ (already outperforming the compared methods) and $N=20,50,100$, ideally, LLR-BC incurs only an additional computational overhead of less than 20%, 8%, and 4%, respectively, compared with fine-tuning. Under our implementation, for all experiments, LLR-BC introduces no more than 8 additional hours based on fine-tuning of training for the whole lifelong learning process. Given its superior performance, this additional training complexity is a reasonable and acceptable trade-off.
@@ C.1 Baselines
To adapt Li (inter), Li (intra), and Feng to our focused scenarios where the generation of new problem instances is uncontrollable, we maintain one instance buffer for each seen task instead of generating new instances for their experience replay, so that each buffer contains the instances generated during learning the corresponding task. Buffers are updated with reservoir sampling. The buffer size of each task is 200, so that there is a total of 1000 buffered instances that can be used for learning the 6th tasks, aligning with the buffer size of LLR-BC. We set $\alpha=0.5$ arbitrarily for Li (inter), Li (intra) and Feng, as we did not find suggested value or method for setting the value of $\alpha$.. Feng contains a learnable model parameter $\mathbf{B}$ with fixed size $n\times n$ ($n$ is the problem scale). We did not find the description about the way to handle $\mathbf{B}$ under scenarios where problem scale changes, so we removed $\mathbf{B}$ to make it applicable to our scenarios. To adapt LiBOG in VRP tasks, we directly embed its two key modules, i.e., the intra-task and inter-task consolidation terms, into POMO for lifelong learning. We set weight of the consolidation term to $10$ for EWC and LiBOG, and the weight of the intra-task consolidation term to 1e41e-4 for LiBOG. For all runs, the solver learns the first task from an identical randomly generated model. As Restart learns each task independently under any task order without being affected by the order, we run Restart once on each task and calculate metrics under different task orders. Therefore, Restart’s APl owns not std. value. As in most cases, POMO can complete the learning on each single task for 200 epochs (Kwon et al., 2020), we let each method train for 200 epochs on each task. To ensure comparable training time across tasks of different problem scales, we vary the number of training problems sampled per epoch: 10,000 for tasks of scale 20, 4,000 for scale 50, and 2,000 for scale 100, with batch sizes of 64, 32, and 16, respectively. Following  Kwon et al. (2020), during training, we sample a batch of problem instances and run $N$ parallel solving processes per instance, each corresponding to a different starting node, where $N$ is the problem scale. This results in $N^{2}$ parallel trajectories per batch, yielding $N^{2}$ independent states and behaviors per experience batch. The size of the experience batch varies across tasks with different problem scales. We consider this reasonable, as larger-scale tasks are typically more complex and thus require larger experience batches to effectively retain learned knowledge. We experimented with buffering and sampling at the level of individual experiences rather than experience batches. However, when using GPU-based parallelization, operating at the batch level significantly reduces computation time without yielding significant performance differences.
@@ C.2 Evaluation Metrics
C.3 Ablation Study Details
In the ablation version of the confidence measure, negative entropy of action probability distribution, and Top-2 margin, i.e., the absolute difference between top-2 action probability, are used instead of variance. We use the maximal possible entropy to normalize entropy, similar to the normalization of variance. Top-2 margin has a value range of [0,1], so it does not need normalization. All three measurements, including variance, intend to assign higher confidence values (thus lower weights) when the model selects the estimated best action with higher probability, as the consolidation process requires. Results show that, although T2M performs slightly worse in terms of forgetting (arguably because it focuses only on the top two actions and thus discards information about the others), the overall performance of the three measures is very similar. As negative entropy is commonly used as a confidence measurement in recent works, we further calculate the coorelation between varaince and negative entropy. We compute both the variance and the negative entropy for all experiences stored in the buffer after running LLR-BC once under task order 1. The Pearson correlation coefficient between the two measurements is 0.972367 (where 1 indicates perfect positive correlation), showing that, in practice, variance tracks confidence during the lifelong learning process very well. This suggests that, while variance is a simple and effective choice, LLR-BC is not very sensitive to the specific confidence measure. Since LLR-BC is designed to assign higher weights to lower-confidence steps, other confidence measures with similar properties can also be used.
For instance-based buffering, we conduct batch-level buffering, set the buffer size to 20, so that the expected buffered steps are more than the buffer size in the step-based buffer (20 * $\frac{20+20+50+50+100+100}{6}\approx 1133\textgreater 1000$). And 1 (batch of) instances will be sampled from the buffer for each consolidation, i.e., $\mathcal{E}=1$. With the similar buffer size (in terms of buffered steps (state-behavior pairs) ), instance-based design reduces the instance-diversity of the buffer, potentially leading to inferior performance. The experimental results indicate that our state-based buffer design outperforms the instance-based buffer design, supporting the its advantage.
Though reservoir sampling gives each experience an equal probability of being kept in the buffer, since the learning processes of different tasks generate different numbers of experiences, the experiences in the buffer are not uniformly added from all tasks. For tasks with a larger problem scale, more experiences from these tasks are buffered. For one run on task order 1 with buffer size 1000, after learning all tasks, the numbers of experiences in the buffer from each task are: E:162, C:255, G:141, U:78, R:77, GM:287. The ratio of the numbers of buffered experiences is highly correlated with the ratio of the problem scales. We suggest this is reasonable, as tasks with larger scale may be more difficult and need more experiences for consolidating their knowledge and addressing catastrophic forgetting during later learning. After training all tasks with the rescaling version -Res, the numbers of experiences in the buffer from each task are: E:148, C:148, G:144, U:199, R:208, GM:153. Though the buffer is more balanced, it performs (cf. Table 3) very similarly to (or even slightly worse than) the original LLR-BC, suggesting the effectiveness of our design.
C.4 Other Details
@@ D.1 Main Experiments on CVRP
The average solution distance and optimality gap (optimal solutions are obtained by HGS) of the solver trained with each method on all tasks follwoing task order 1 are listed in Table 6. Notably, we use the solver trained after the corresponding task to evaluate Restart, making it single task method.
@@
Order Method k=3 k=6 AP AF AMF APl AG AP AF AMF APl AG 1 Li (inter) 20.7 0.0 0.0 21.2 26.3 24.3 0.1 0.1 24.8 27.7 Li (intra) 33.4 0.0 0.0 39.4 47.8 33.2 0.0 0.0 37.2 41.5 Feng 23.2 1.4 1.4 18.8 27.6 25.7 2.2 2.9 23.9 34.4 Restart 72.0 8.2 8.2 11.8 36.3 95.7 27.1 27.7 9.1 40.2 Fine-tuning 32.8 3.2 3.2 7.5 32.8 34.5 19.9 23.1 4.6 31.3 EWC 35.9 3.1 3.1 8.6 32.6 39.6 19.7 21.5 7.5 33.5 LiBOG 38.3 2.7 2.7 8.9 33.4 47.7 18.0 18.6 8.3 34.7 LLR-BC 8.0 0.0 0.0 7.0 27.1 4.7 0.6 0.7 4.1 23.3 2 Li (inter) 28.6 0.1 0.1 30.4 47.2 33.2 0.0 0.0 34.4 42.2 Li (intra) 33.1 0.5 0.5 41.1 59.9 37.4 0.0 0.2 41.8 51.6 Feng 27.3 13.6 13.6 26.2 64.7 25.3 3.4 5.8 24.5 43.2 Restart 44.9 92.5 92.5 8.7 60.7 42.1 37.2 64.4 9.1 52.6 Fine-tuning 21.8 29.9 29.9 3.7 50.0 18.8 14.3 23.3 2.5 38.4 EWC 19.5 33.1 33.1 4.4 48.6 19.7 12.4 23.3 4.9 37.9 LiBOG 22.4 29.1 29.1 4.9 51.1 22.2 12.1 21.2 5.4 39.2 LLR-BC 2.8 0.0 0.0 3.3 37.1 2.2 0.3 0.3 1.8 22.6 3 Li (inter) 20.2 0.0 0.0 21.3 26.8 24.0 0.0 0.0 24.8 28.2 Li (intra) 30.3 0.0 0.0 36.5 44.8 35.6 0.0 0.0 39.2 44.5 Feng 19.1 2.9 2.9 15.0 21.5 25.6 1.3 2.4 24.2 26.6 Restart 106.8 10.7 10.7 7.8 30.9 95.7 22.8 35.7 9.1 51.9 Fine-tuning 35.6 4.9 4.9 6.1 24.1 33.4 21 27.4 3.9 44.4 EWC 40.3 4.1 4.1 6.7 22.3 38.0 18.0 25.4 6.8 40.5 LiBOG 42.6 4.3 4.3 6.8 25.0 43.0 18.4 25.9 6.7 41.7 LLR-BC 6.4 0.0 0.0 5.9 23.0 5.0 0.9 1.0 3.7 23.1 4 Li (inter) 36.5 0.2 0.2 40.9 62.2 40.1 0.1 0.1 42.8 52.5 Li (intra) 31.7 0.3 0.3 44.8 75.9 32.5 0.0 0.1 40.3 54.8 Feng 22.1 10.6 10.6 20.9 66.3 20.7 2.1 4.3 25.4 52.4 Restart 35.7 44.5 44.5 10.3 77.3 42.1 26.5 35.5 9.1 52.1 Fine-tuning 14.6 32.6 32.6 7.0 70.0 20.2 18.9 29.9 4.5 45.8 EWC 19.2 28.4 28.4 8.4 69.6 27.1 18.5 23.6 7.9 46.4 LiBOG 19.4 29.1 29.1 8.7 69.0 26.5 20.0 24.3 7.9 46.7 LLR-BC 7.0 0.3 0.3 6.4 58.2 4.7 0.3 0.6 4.0 32.6 5 Li (inter) 29.4 0.1 0.1 31.7 45.1 38.6 0.0 0.0 41.2 50.1 Li (intra) 28.5 0.1 0.1 36.7 53.3 31.9 0.0 0.1 37.9 48.4 Feng 26.3 1.3 1.3 24.5 47.6 25.9 7.0 7.0 23.2 38.9 Restart 27.2 19.3 19.3 9.5 58.6 26.7 93.1 96.5 9.1 50.5 Fine-tuning 11.8 11.8 11.8 7.0 50.2 10.5 25.5 36.7 3.7 50.6 EWC 17.8 14.5 14.5 8.2 49.1 17.0 29.0 33.7 7.5 40.9 LiBOG 18.3 13.5 13.5 8.6 49.0 17.3 29.9 35.7 7.7 41.7 LLR-BC 6.5 0.4 0.4 6.7 45.3 4.2 1.6 1.6 3.7 31.8
Table 6: Avg. solution distance (optimality gap) of each method on CVRP.
Method E C G U R GM Li (inter) 10.74 (3.5%) 15.42 (6.1%) 11.59 (5.2%) 6.73 (4.1%) 5.83 (19.1%) 13.44 (6.6%) Li (intra) 10.89 (5.0%) 15.61 (7.3%) 11.71 (6.4%) 6.76 (4.5%) 5.85 (19.4%) 13.53 (7.3%) Feng 10.81 (4.2%) 15.54 (6.9%) 11.52 (4.5%) 6.72 (3.9%) 5.79 (18.1%) 13.59 (7.7%) Restart 10.74 (3.5%) 15.28 (5.1%) 11.38 (3.3%) 6.57 (1.6%) 5.71 (14.4%) 13.17 (4.3%) Fine-tuning 10.93 (5.4%) 15.40 (5.9%) 11.80 (7.1%) 7.04 (8.9%) 5.87 (17.6%) 13.03 (3.3%) EWC 10.97 (5.7%) 15.41 (6.0%) 11.86 (7.6%) 7.04 (8.9%) 5.97 (19.5%) 13.10 (3.8%) LiBOG 11.00 (6.0%) 15.43 (6.1%) 11.89 (7.9%) 7.27 (12.3%) 6.01 (20.3%) 13.10 (3.8%) INViT 11.18 (7.8%) 15.96 (9.8%) 11.73 (6.5%) 7.03 (8.6%) 5.97 (19.7%) 13.77 (9.2%) POMO-MT-EL 10.66 (2.7%) 15.10 (3.9%) 11.38 (3.3%) 6.59 (1.8%) 5.71 (14.5%) 13.04 (3.3%) POMO-MT-BL 10.66 (2.8%) 15.10 (3.9%) 11.36 (3.1%) 6.58 (1.7%) 5.71 (14.5%) 13.04 (3.4%) LLR-BC 10.72 (3.3%) 15.17 (4.3%) 11.36 (3.1%) 6.56 (1.5%) 5.70 (14.1%) 13.03 (3.3%)
Table 7: Avg. solution distance (optimality gap) of each method on TSP.
Method E C G U R GM Li (inter) 5.57 (1.4%) 6.35 (4.8%) 5.59 (2.5%) 3.84 (0.4%) 1.98 (0.2%) 4.47 (8.8%) Li (intra) 5.76 (4.8%) 6.59 (8.9%) 5.60 (2.8%) 3.90 (2.0%) 2.00 (1.4%) 4.69 (14.0%) Feng 5.66 (2.9%) 6.40 (5.7%) 5.58 (2.4%) 3.86 (1.0%) 1.98 (0.2%) 4.51 (9.7%) Restart 5.57 (1.3%) 6.31 (4.3%) 5.49 (0.5%) 3.83 (0.1%) 1.99 (0.8%) 4.25 (3.3%) Fine-tuning 5.67 (3.2%) 6.32 (4.4%) 5.72 (4.8%) 3.88 (1.4%) 1.98 (0.3%) 4.20 (2.1%) EWC 5.67 (3.3%) 6.36 (5.0%) 5.61 (2.7%) 3.89 (1.6%) 1.98 (0.4%) 4.23 (2.8%) LiBOG 5.66 (3.1%) 6.33 (4.5%) 5.71 (4.6%) 3.88 (1.5%) 1.98 (0.4%) 4.23 (2.9%) INViT 5.75 (4.6%) 6.53 (7.9%) 5.66 (3.8%) 3.93 (2.8%) 2.00 (1.4%) 4.58 (11.4%) POMO-MT-EL 5.61 (2.1%) 6.25 (3.2%) 5.48 (0.3%) 3.86 (0.8%) 1.98 (0.1%) 4.25 (3.3%) POMO-MT-BL 5.59 (1.8%) 6.24 (3.0%) 5.47 (0.3%) 3.85 (0.6%) 1.98 (0.1%) 4.24 (3.1%) LLR-BC 5.56 (1.3%) 6.21 (2.6%) 5.46 (0.1%) 3.83 (0.2%) 1.98 (0.0%) 4.20 (2.1%)